Given: You put away S10,000 at the end of each year for an anticipated future worth of $74420 in 5 years. Find: The annual interest rate. Problem 2.6 Annual deposits earning monthly interest Given: You deposit $10,000 each year at a monthly interest of 8%. Find: The amount you have in 10 years.
Answer #1

Problem 2.5

yearly payment or PMT=10000

future value=74420

N=5

FV=PMT*(1+r)^4+PMT*(1+r)^3+PMT*(1+r)^2+PMT*(1+r)^1+PMT*(1+r)^0

74420=10000*(1+r)^4+10000*(1+r)^3+10000*(1+r)^2+10000*(1+r)^1+10000*(1+r)^0

use trial and error method to find value of r , so that both sides of the equation equal.

trying many values of r, we get r=20.00%, this is the annual interest rate

Problem 2.6

annual effective interest rate=(1+(8%/12))^12-1=8.30%

use future value of annuity method

FV or amount in 10 years=P*(((1+r)^n-1)/r)

=10000*(((1+8.30%)^10-1)/8.30%)

=146945.82

the above is answer..